skip to content
Linda
User Tools
Log In
Site Tools
Search
Tools
Show pagesource
Old revisions
Backlinks
Recent Changes
Media Manager
Sitemap
Log In
>
Recent Changes
Media Manager
Sitemap
Trace:
•
procedures
Sidebar
Predmety
Seminar KFM
Python
(Linux ve fyz. lab.)
Zariadenia
Bizhub
Siet/pocitace
TEM
AFM
Rezisto
webcams
Users
jove
michal
tem:procedures
Table of Contents
Procedures
Start work
End work
Alignment
After bakeout
Beam shower
After power failure
Normal vacuum levels
Procedures
Start work
HT→FEG Standby Mode→Normal (it takes ~13min)
Mount sample
Check nitrogen pressure!
Insert holder till the first notch Air→Pump (wait for green light)
Fill ACD with liquid nitrogen (in two steps)
Fully insert holder
Select holder type!
End work
Stage Neutral
Check nitrogen pressure!
Pull holder till last notch
Pump→Air, wait for nitrogen and remove the holder
Mount ACD heater, Maintenance→ACD & Bake→ACD Heat→ON (turns off automatically)
HT→FEG Standby Mode→Standby
Alignment
GUNA(F3) Spot1 center with SHIFT ↔ BRIGHT TILT Spot5 center with SHIFT
GUNA(F3) and Anode wobler(F5) make symetric with DEF
COND STIG: make round with DEF
STD FOCUS adjust Z (minimal contrast, no hallo, or no motion with IMAGE WOBB)
Tilt and shift purity (at correct focus, and diffraction settings!)
OBJ wobler (F4) make feature stationary with DEF(BRIGHT TILT)
make fft round with DEF(OBJ STIG)
SPECTRUM→move ZLP to center using FL lens
Filter→Slit→IN move slit to center using position
back to IMAGE, center circle with Isochromacity
center diffraction with DEF(PLA)
focus diffraction with DIFF FOCUS at Alpha 3, camera length 200cm, condensor fully CW
After bakeout
remove mechanical safeties from EDX detector valve and beam valve
Cycle BEAM VALVE with HT OFF
Open EDX detector valve with BEAM VALVE closed
Conditioning knob → COND
HT→HT Conditioning→Auto Procedure→Start (it takes ~60 min)
Conditioning knob → OPERATE
Beam shower
Camera length 150cm
CLA out, SAA #4, HCA #4
CL1 0x8000, CL2 0x6000 (load STEMBeamShower.flc in FLC Panel)
wait 10-15 min
After power failure
reset chiller complaining about closed outlet
(on-off HT breaker at the back of power supply unit)
press (longer) “I” on the front of the microscope
(Lenses on - front of the microscope)
(EDX detector controller on-off, while SIP is not running (ACD on))
watch vacuum
(HT conditioning: switch knob to Cond, Auto procedure start, switch knob back to Operate)
HT → Auto HT & Emission → Start up
Normal vacuum levels
Column/PIG1
~25 uA
SIP starts at ? uA
Specimen/PIG4
~25 uA
~250 uA without holder, pump LED goes green at 28 uA
Camera/PIG3
~25 uA
RT1/PIG5
~80 uA
RT2/PIG6
~25 uA
20l SIP
<1 uA
60l SIP
~1 uA
Column SIP
200 uA
800 uA with specimen without LN2 in ACD
tem/procedures.txt
· Last modified: 2021/10/25 08:53 (external edit)
Page Tools
Show pagesource
Old revisions
Backlinks
Back to top